OpenStudy (anonymous):

Ready to pull my hair out on this one! I can get the first half correct..... Find any x intercepts on the graph of the equation: y=5x^2+7x-5

OpenStudy (anonymous):

5x^2+7x-5=0 a=5 b=7 c=-5

OpenStudy (anonymous):

Right?

jimthompson5910 (jim_thompson5910):

so far so good

jimthompson5910 (jim_thompson5910):

now use the quadratic formula

OpenStudy (anonymous):

Im not sure hot to put in the correct symbols on this screen

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(7)\pm\sqrt{(7)^2-4(5)(-5)}}{2(5)}\] \[\Large x = \frac{-7\pm\sqrt{49-(-100)}}{10}\] \[\Large x = \frac{-7\pm\sqrt{49+100}}{10}\] Keep going...

OpenStudy (anonymous):

Yes! I got this far after about an hour.... I will write it out

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):